# Deep Joint Source Channel Coding With Attention Modules Over MIMO Channels

### 12 Mar 2024 v2

@@ IV-B Entropy of DJSCC encoder output over sub-channels
The goal of non-parameter entropy estimation is to estimate the differential entropy of the underlying distribution only from a finite number of samples when the probability density distribution is unknown[19]. The mainstream methods of non-parameter entropy estimation include kernel density estimators(KDEs) and $k$ nearest neighbors (kNNs), etc. In our experiment, we adopt the kNN entropy estimation method, whose error rate is no more than twice of the Bayesian error rate (minimum error)[20]. The core idea of kNN methods is to estimate the probability mass of local field around each sample by making Gaussian approximation for $k$ nearest points around the sample[19]. The calculation formula of estimated differential entropy $\widehat{H}(\boldsymbol{X})$ is given by
@@
where $\psi$ denotes the digamma function, $N$ denotes the total number of samples, $k$ denotes the number of samples in the local field, $c_{d}$ is a constant term that denotes the unit volume of the local field, $d$ denotes the dimension of the data, and $\varepsilon(i)$ denotes the distance from the $i$th sample to the $k$th nearest neighbor point. Details of the formula and derivation process can be found in[19, 21]. In our simulation, we set $d=2$ for transmitting complex symbols and $k=\sqrt{N}$ without loss of generality[22].
In order to observe how differential entropy changes with different sub-channel qualities, we generate MIMO channel matrices with fixed singular values , which satisfy the power constraint $\sum\lambda_{i}^{2}=C$. We consider $2\times 2$ MIMO channel and the SNR of 10 dB. The estimated entropy for different singular value ratios are shown in Fig. 6. When the sub-channels’ singular value ratio $r=1$, the two sub-channels have the same equivalent SNR as well as channel capacity, and the data streams going through the two sub-channels also have very close differential entropy values. As the channel singular value ratio grows, the capacity of one sub-channel rises while the capacity of the other sub-channel falls. We can observe that the differential entropy of the encoder output follows the same pattern: the higher the channel capacity, the higher the differential entropy, and vice versa. Meanwhile, it can be clearly seen that the differential entropy difference between the two sub-channels is also increasing. By exploiting the entropy estimation, we reveal the mechanism of DJSCC in different sub-channels, i.e., more information is transmitted in the better sub-channel.
@@ V Conclusion
In this paper, we design two DJSCC structures with attention modules for the MIMO channel, including a serial structure and a parallel structure. We apply the new MIMO-DJSCC schemes for image transmission, and both network structures show improved image recovery performance compared to the existing DJSCC network. Furthermore, we investigate the estimated entropy of MIMO-DJSCC encoder over different sub-channels and observe that more information is transmitted in the MIMO subchannel of a higher SNR. The combination of attention modules and SVD-based precoding enables the encoding network to adapt to the CSI of different sub-channels, generate data streams of different entropy. Thus, the MIMO-DJSCC could exploit the spatial gain of the MIMO channel, that results in improved performance than the SISO channel.

### 15 Mar 2024 v3

@@ IV-B Entropy of DJSCC encoder output over sub-channels
The goal of non-parameter entropy estimation is to estimate the differential entropy of the underlying distribution only from a finite number of samples[23]. In our experiment, we adopt the kNN entropy estimation method, whose error rate is no more than twice of the Bayesian error rate (minimum error). The core idea of kNN methods is to estimate the probability mass of local field with $k$ nearest points around the sample[23]. The calculation formula of estimated differential entropy $\widehat{H}(\boldsymbol{X})$ is given by
@@
where $\psi$ denotes the digamma function, $N$ denotes the total number of samples, $k$ denotes the number of samples in the local field, $c_{d}$ is a constant term that denotes the unit volume of the local field, $d$ denotes the dimension of the data, and $\varepsilon(i)$ denotes the distance from the $i$th sample to the $k$th nearest neighbor point. Details of the formula and derivation process can be found in[23]. In our simulation, we set $d=2$ and $k=\sqrt{N}$ without loss of generality.
In order to observe how differential entropy changes with different sub-channel qualities, we generate MIMO channel matrices with fixed singular values , which satisfy the power constraint $\sum\lambda_{i}^{2}=C$. We consider $2\times 2$ MIMO channel and the SNR of 10 dB. The estimated entropy are shown in Fig. 6. When the sub-channels’ singular value ratio $r=1$, the two sub-channels have the same equivalent SNR as well as channel capacity, and the data streams going through the two sub-channels also have very close differential entropy values. We can observe that the differential entropy of the encoder output follows the pattern: the higher the channel capacity, the higher the differential entropy, and vice versa. By exploiting the entropy estimation, we reveal the mechanism of DJSCC in different sub-channels, i.e., more information is transmitted in the better sub-channel.
@@ V Conclusion
In this paper, we design two DJSCC structures with attention modules for the MIMO channel, including a serial structure and a parallel structure. We apply the MIMO-DJSCC schemes for image transmission, and both network structures show improved image recovery performance. Furthermore, we investigate the estimated entropy of MIMO-DJSCC encoder over different sub-channels and observe that more information is transmitted in the better sub-channel. The combination of attention modules and SVD-based precoding enables the network to adapt to different sub-channels, generate data streams of different entropy. Thus, the MIMO-DJSCC could exploit the spatial gain of the MIMO channel, that results in improved performance than the SISO channel.